OpenStudy (anonymous):

one number plus twice another number, gives a sum of 42, the difference of these same two numbers equals -3. *what are the two numbers?

OpenStudy (anonymous):

a+2b = 42 a-b = -3 ---> a = -3+b (-3+b) +2b = 42 ----> b= 15 a= -3 + (15) = 12 Check: 12+ 2(15) = 42 ? Yes 12 - 15 = -3 ? yes

OpenStudy (anonymous):

so what are the two numbers?

OpenStudy (anonymous):

If you do the work shown above and actually read the solution then you should be able to figure it out... especially since i wrote the answer up there...

OpenStudy (anonymous):

ty